OpenStudy (anonymous):

which of the following forms a proportion of 24/18? A.18/12 B.20/15 C.3/4 D.14/8

OpenStudy (anonymous):

one way to do this is to reduce \[\frac{24}{18}=\frac{4}{3}\]

OpenStudy (anonymous):

but since none of your answers are \(\frac{4}{3}\)we can see if we can find an equivalent fraction to it

OpenStudy (anonymous):

that is not too hard \[\frac{4}{3}=\frac{4}{3}\times \frac{5}{5}=\frac{20}{15}\]

OpenStudy (anonymous):

so it would be b

OpenStudy (anonymous):

yes it would would you like to check it is right?

OpenStudy (anonymous):

or let me ask you this do you know how to check it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[\frac{24}{18}=\frac{20}{15}\]you can check if \[24\times 15=20\times 18\] or not

OpenStudy (anonymous):

since both times you get \(360\) then the fractions must be equal

OpenStudy (anonymous):

thanks